if qx=25-2.5px+5/3py+0.5I where px=20 py=15 and I=100. what is the cross elasticity of demand
1
Expert's answer
2018-11-13T12:19:09-0500

if qx=25 - 2.5px + 5/3py + 0.5I, px = 20, py = 15 and I = 100, then qx = 25 - 2.5*20 + 5/3*15 + 0.5*100 = 50 units. The cross elasticity of demand is:

Ed = -b*py/qx = 5/3*15/50 = 0.5, so these goods are substitutes.